# CineTrans: Learning to Generate Videos with Cinematic Transitions via Masked Diffusion Models

### v1

@@ 6.2 Evaluation
To comprehensively evaluate the multi‑shot video generation task, we design 100 prompts spanning four categories using GPT‑4o (Achiam et al., 2023), each incorporating cinematic transitions. For each initial general description, we manually annotate the shot count and then generate captions for each shot, thus constructing a complete prompt series. During evaluation, each method employs its required prompt format (multi‑prompt or single‑prompt) for sampling. To ensure a fair comparison, when using only the general description, the shot count is explicitly included in the text.
Metrics. We propose a set of metrics to evaluate the generated videos from three perspectives: transition control, temporal consistency, and overall video quality. First, we evaluate transition control by performing shot segmentation (Soucek & Lokoc, 2024), comparing the detected number of shots in the generated video against the prompt’s specified count, and computing the Transition Control Score as detailed in the Appendix E. Second, temporal consistency encompasses both intra-shot and inter-shot consistency. For intra-shot coherence, following VBench (Huang et al., 2024), we treat each shot as a video and compute both subject and background consistency, averaging the results across all shots. For inter-shot coherence, we evaluate from both semantic and visual perspectives. Semantically, we extract features for each shot using ViCLIP (Wang et al., 2023b) and compute cosine similarities. Visually, following VBench-Long, we compute the similarity between the middle frames of each shot to derive both subject and background similarity scores, which are then averaged to obtain the final inter-shot visual consistency. Furthermore, as discussed in Section 3, inter-shot consistency prioritizes high-level coherence aligned with film editing conventions, rather than low-level pixel similarity. Therefore, a high inter-shot consistency score, approaching the theoretical maximum, may actually indicate high pixel-level similarity, which violates the prior of multi-shot video design. To address this limitation, we introduce an auxiliary metric: Consistency Gap with Film Editing. For inter-shot consistency metrics, we collect reference score distributions from 1000 randomly selected multi-shot videos that have undergone professional film editing. The Jensen-Shannon Distance between the score distribution of generated videos and that of the reference set is computed as the Consistency Gap. Figure 7(a) illustrates the convergence of the consistency distribution for our selected reference set. This metric complements the raw consistency score by quantifying the deviation from natural film-editing style, thus providing a more comprehensive evaluation of video consistency. Finally, following VBench, we assess overall video quality using aesthetic quality and overall consistency, which also evaluate the visual and semantic aspects, respectively.
@@
Results. The quantitative comparison is shown in Table 3. Even in the training‑free setting, our proposed mask mechanism delivers strong transition control. In terms of consistency, CineTrans achieves high scores and closely aligns with film‑editing style, which is also shown in Figure 7(b). Because aesthetic quality is largely determined by the base model, CineTrans‑Unet performs slightly worse in this regard, whereas CineTrans‑DiT exhibits superior results. Furthermore, as shown in Table 4, the user study results demonstrate the superiority of our method in terms of user preference.
@@
Table 4: Results of User Study
Model Transition Control Consistency
CineTrans-DiT (ours) 4.60 $\pm$ 0.50 4.15 $\pm$ 0.75
CineTrans-Unet (ours) 4.75 $\pm$ 0.44 4.10 $\pm$ 0.72
StoryDiffusion+CogVideoXI2V - 3.95 $\pm$ 0.76
HunyuanVideo+Cinematron 3.60 $\pm$ 1.95 3.80 $\pm$ 0.68
HunyuanVideo 3.45 $\pm$ 1.88 3.50 $\pm$ 0.76
CogVideoX 2.50 $\pm$ 1.24 3.05 $\pm$ 0.60
Wanx2.1-T2V-turbo 3.25 $\pm$ 1.77 3.45 $\pm$ 0.69
@@ 6.3 Ablation Studies
We conduct ablation studies to assess the impact of the mask mechanism and fine-tuning process. The results shown in Table 2 demonstrate that our proposed components effectively enhance performance. In terms of inter‑shot visual consistency, the fine‑tuned model yields a lower Consistency Score, reflecting increased compositional variation between shots introduced by training on film‑edited multi‑shot videos. This effect is further corroborated by a reduced Consistency Gap, indicating closer alignment with the film‑editing style. The slight decline in Aesthetic Quality after fine‑tuning may be attributed to aesthetic domain differences between Cine250K and the original training set of base model. Furthermore, Figure 8 illustrates how fine‑tuning and Visible-First‑Frame Attention enhance video consistency and enable stable generation.
@@ A.1 Method of Splitting and Shot Segementation
During the Splitting stage, videos are initially segmented using PySceneDetect (Castellano, 2024). In the subsequent gradual-change removal stage, TransNetV2 (Soucek & Lokoc, 2024) is employed to detect and handle gradual transitions, yielding the final shot boundaries and associated shot labels. This pipeline is selected because Pyscenedetect operates exclusively on the CPU, offering greater processing efficiency than Transnetv2; however, for a pre-segmented video, Transnetv2 achieves higher shot segmentation accuracy and can handle gradual transitions more effectively than Pyscenedetect.
@@ A.2 Details of Stitching Phase
During the Stitching stage, adjacent segments exhibiting high semantic similarity are merged to form a single video, thereby constructing multi-shot video collections. Specifically, we extract semantic features from the first and last frames of each segment using ImageBind (Girdhar et al., 2023) and quantify their similarity via Euclidean distance. For the $i$-th segment $C^{i}$, we denote the semantic features of its first and last frames as $C_{\mathrm{first}}^{i}$ and $C_{\mathrm{end}}^{i}$, respectively, with their distance represented as $\mathrm{dis}(C_{\mathrm{first}}^{i},C_{\mathrm{end}}^{i})$. Based on the distance, video segments are processed sequentially and either stitched or filtered according to the following criteria.
@@ Appendix B Statistic of Cine250K
As we present in Section 4, Cine250K is a carefully curated multi-shot video dataset with detailed captions. This section presents its overall statistics. As shown in Figure 11, the average video duration is 10.75s, and the average caption length is 148.79. Most videos contain 2 to 3 shots. It is important to note that although duration and shot count filtering are applied during data processing, TransnetV2 (Soucek & Lokoc, 2024) is later utilized to remove gradual changes and re-identify shots. As a result, the final shot count and duration do not strictly adhere to the initial filtering criteria. After reidentification, 87.99% of the videos contain 2 to 5 shots. Figure 11 presents the shot distribution for videos with 1 to 10 shots, which represents 99.90% of all videos.
@@
In Figure 12, we select several example video-text pairs to demonstrate the specific characteristics of the cinematic multi-shot sequences in the dataset and the style of captions.
@@ C.2 Multi-prompt
In Section 6.2, we note that some methods employ a multi-prompt strategy at inference, i.e., each shot is prompted by its own text description. CineTrans-DiT also supports this capability. In addition to the primary mask matrix between video tokens, we introduce an additional mask between text and video tokens, following Qi et al. (2025), to enable precise semantic control over each shot. It is worth noting that this mask is applied only to the attention layers governing text–video token interactions, and is distinct from our proposed mask mechanism, which operates on video–video token interactions.
@@ Appendix D Additional Results
As for the architecture, we investigate both the temporal-spatial-decoupled framework and the full attention framework. The temporal-spatial-decoupled framework applies the temporal attention module directly to the time sequences, where tokens at different spatial locations do not interact with each other. In contrast, the full attention framework operates on all tokens of the video, allowing correlations across both temporal and spatial dimensions simultaneously. To focus on frame correlations, we group and average the tokens in the full attention framework by the frames before conducting further analysis. In Figure 14 and Figure 15, we use Wang et al. (2024) and Kong et al. (2024) as representatives of different architectures and present the attention maps across different layers and timesteps in both qualitative and quantitative manners.

### v2

@@ 5.2 Evaluation
For comprehensive evaluation, we design 100 hierarchical prompts with transitions using GPT‑4o (achiam2023gpt), where each initial general description is annotated with shot count and then expanded into shot-level captions, thus constructing a complete prompt. During evaluation, each method uses its required prompt format (multi‑prompt or single‑prompt). To ensure fair comparison, when using only the single general prompt, the shot count is explicitly included in the text.
Metrics. We evaluate generated videos along three dimensions: transition control, temporal consistency, and overall quality. For transition control, we assess whether the shot count in the generated video matches the specified count using the Transition Control Score, computed as in Equation 5, where $s_{\mathrm{generated}}$ and $s_{\mathrm{specified}}$ denote the shot counts in the generated video and the prompt, respectively, and $k$ is a hyperparameter controlling the tolerance margin.
$x=\dfrac{s_{\mathrm{generated}}}{s_{\mathrm{specified}}}$ (4)
$\mathrm{Transition\,Control\,Score}=\dfrac{x^{k}}{e^{k(x-1)}}$ (5)
Second, temporal consistency includes both intra- and inter-shot coherence. For intra-shot, following VBench (huang2024vbench), we compute subject and background consistency for each shot and average the results. For inter-shot, we consider both semantics and visuals. Semantically, we compute cosine similarity of ViCLIP (internvid) features across shots. Visually, following VBench-Long, we compute subject and background similarity between middle frame of shots and average them. Furthermore, as discussed in Section 4.1, overly high inter-shot scores may reflect undesirable pixel-level similarity, which violates the prior of multi-shot video design. To address this limitation, we introduce Consistency Gap as an auxiliary metric, defined as the Jensen-Shannon Distance between the score distribution of generated videos and that of the reference set:
$\mathrm{JSD}(P,Q)=\sqrt{\frac{1}{2}D\Big(P\parallel\frac{P+Q}{2}\Big)+\frac{1}{2}D\Big(Q\parallel\frac{P+Q}{2}\Big)}.$ (6)
Figure 7(a) illustrates the convergence of the consistency distribution for the reference set, which consists of 1,000 professionally edited multi-shot videos. This metric complements the raw consistency score by quantifying the deviation from natural film-editing style, thus providing a more comprehensive evaluation of video consistency. Finally, following VBench, we assess overall video quality using aesthetic quality and overall consistency, which also evaluate the visual and semantic aspects, respectively.
@@
Results. The quantitative comparison is shown in Table 1. Our proposed CineTrans achieves near-perfect cinematic transition control across all prompts. In terms of consistency, CineTrans achieves high scores and closely aligns with film‑editing style, which is also shown in Figure 7(b). Because aesthetic quality is largely determined by the base model, CineTrans‑Unet performs slightly worse in this regard, whereas CineTrans‑DiT exhibits superior results.
@@ 5.3 Ablation Studies
We conduct ablation studies to assess the impact of the mask mechanism and fine-tuning process. The results in Table 2 demonstrate the effectiveness of our components and the reasonable performance of the training-free variant. In terms of inter‑shot visual consistency, the fine‑tuned model yields a lower Consistency Score, reflecting increased compositional variation between shots introduced by training on film‑edited multi‑shot videos. This effect is further corroborated by a reduced Consistency Gap, indicating closer alignment with the film‑editing style. The slight decline in Aesthetic Quality after fine‑tuning may be attributed to aesthetic domain differences between Cine250K and the original training set of base model. Furthermore, Figure 8 illustrates how fine‑tuning and Visible-First‑Frame Attention enhance video consistency and enable stable generation.
Figure 9: Smoother transitions achieved through soft masking.
Additionally, since hard cuts represent the majority of cinematic transitions (99.58%), we implement shot transitions using hard masks. However, there remains a need to explore soft mask strategies to achieve smoother multi-shot video generation, including seamless transitions between frames without visual breaks, as well as fade-in/fade-out effects. We present preliminary results in Figure 9, with further discussion in Appendix D.5. Soft masks improve transition smoothness and consistency but reduce shot transition control, introducing a gap compared to cinematic transitions. Thus, hard masks remain the primary method for shot transitions. Nonetheless, the promising results of the soft mask approach highlight CineTrans’s flexibility and robustness, demonstrating its ability to adapt to a wider range of transition styles.
@@ 6 Conclusion
Acknowledgment
The work was supported by Shanghai Artificial Intelligence Laboratory PJ-PRJ24ARC001-20082025. We gratefully acknowledge the support and resources provided, which made this research possible.
Ethics Statement
This work makes use of video data collected from the publicly available Vimeo platform. We only use videos that are publicly accessible and do not require authentication or bypassing of any access restrictions. The dataset is constructed solely for research purposes, in compliance with Vimeo’s Terms of Service, and will not be used for any commercial applications. To mitigate potential privacy concerns, we avoid videos with personally identifiable information and exclude sensitive content. We will not redistribute the raw videos; instead, we provide scripts that allow others to download the data directly from Vimeo if permitted by the platform.
Reproducibility Statement
We have made significant efforts to ensure the reproducibility of our work. The source code is provided in the supplementary materials. The dataset construction procedure is described in detail in Section 3 and Appendix A, which enables other researchers to replicate the data preprocessing pipeline. Together, these resources are intended to facilitate the reproduction and verification of our experimental results.
@@ A.1 Method of Splitting and Shot Segementation
During the Splitting stage, videos are initially segmented using PySceneDetect (pyscenedetect). In the subsequent gradual-change removal stage, TransNetV2 (soucek2024transnetv2) is employed to detect and handle gradual transitions, yielding the final shot boundaries and associated shot labels. This pipeline is selected because Pyscenedetect operates exclusively on the CPU, offering greater processing efficiency than Transnetv2; however, for a pre-segmented video, Transnetv2 achieves higher shot segmentation accuracy and can handle gradual transitions more effectively than Pyscenedetect.
@@ A.2 Details of Stitching Phase
During the Stitching stage, adjacent segments exhibiting high semantic similarity are merged to form a single video, thereby constructing multi-shot video collections. Specifically, we extract semantic features from the first and last frames of each segment using ImageBind (girdhar2023imagebind) and quantify their similarity via Euclidean distance. For the $i$-th segment $C^{i}$, we denote the semantic features of its first and last frames as $C_{\mathrm{first}}^{i}$ and $C_{\mathrm{end}}^{i}$, respectively, with their distance represented as $\mathrm{dis}(C_{\mathrm{first}}^{i},C_{\mathrm{end}}^{i})$. Based on the distance, video segments are processed sequentially and either stitched or filtered according to the following criteria.
@@ Appendix B Statistic of Cine250K
As we present in Section 3, Cine250K is a carefully curated multi-shot video dataset with detailed captions. This section presents its overall statistics. As shown in Figure 12, the average video duration is 10.75s, and the average caption length is 148.79. Most videos contain 2 to 3 shots. It is important to note that although duration and shot count filtering are applied during data processing, TransnetV2 (soucek2024transnetv2) is later utilized to remove gradual changes and re-identify shots. As a result, the final shot count and duration do not strictly adhere to the initial filtering criteria. After reidentification, 87.99% of the videos contain 2 to 5 shots. Figure 12 presents the shot distribution for videos with 1 to 10 shots, which represents 99.90% of all videos.
Table 5: Comparison of Cine250K and other video-text datasets.
InternVid (internvid) 234M 11.7s 17.6 - High 720P
LVD-2M (lvd) 2M 20.2s 88.7 - Medium Diverse
OpenVid-1M (nan2024openvid) 1M N/A N/A - High Diverse
Panda70M (panda) 70.8M 8.5s 13.2 Medium 720P
LLaVA-Video-178K (zhang2024llavavideo) 178K $\sim$ 40.4s $\sim$ 300 High Diverse
Shot2Story20K (han2023shot2story20k) 20K 16s 201.8 Medium 720P
Shot2Story134K 134K N/A N/A Medium 720P
Cine250K (Ours) 250K 10.75s 148.79 High 720P
@@
In Figure 13, we select several example video-text pairs to demonstrate the specific characteristics of the cinematic multi-shot sequences in the dataset and the style of captions. In Table 5, we compare Cine250K with other datasets. As a multi-shot video dataset with detailed shot labels, the high-quality content of Cine250K can significantly facilitate research and exploration in multi-shot video generation.
@@
Additionally, when Cine250K is used as the training dataset for CineTrans, the distribution of transition point positions and shot lengths is relatively uniform (Figure 14), which enables CineTrans to achieve precise frame-level control without timestamp jitter and ensure stability when varying shot lengths are used as conditions.
Figure 14: Distribution of shot length and transition points in Cine250K as the training dataset. Given the temporal compression of the VAE encoder, temporal slices are used as the unit.
@@ C.2 Multi-prompt
In Section 5.2, we note that some methods employ a multi-prompt strategy at inference, i.e., each shot is prompted by its own text description. CineTrans-DiT also supports this capability. In addition to the primary mask matrix between video tokens, we introduce an additional mask between text and video tokens, following qi2025mask2dit, to enable precise semantic control over each shot. It is worth noting that this mask is applied only to the attention layers governing text–video token interactions, and is distinct from our proposed mask mechanism, which operates on video–video token interactions.
C.3 Transition Point Selection
In Section 4.3, we introduce the Mask Mechanism that achieves temporal control of shot transitions. During evaluation, shot transition points should be predefined in advance to generate multi-shot videos. To address this, we predefine several suitable shot transition points for different shot counts and randomly select them during inference. However, as mentioned in Appendix B, our training set adequately supports shot transitions at various positions, so random sampling of transition points does not lead to issues such as visual collapse (Figure 16). Nevertheless, for optimal visual performance, including action completeness and overall video quality, uniform shot lengths are preferred, as they avoid extremely short shots.
Figure 16: Multi-shot videos with shorter length for last two shots. It can be observed that the visual quality remains intact even when the shot lengths are relatively extreme.
@@ Appendix D Additional Results
As for the architecture, we investigate both the temporal-spatial-decoupled framework and the full attention framework. The temporal-spatial-decoupled framework applies the temporal attention module directly to the time sequences, where tokens at different spatial locations do not interact with each other. In contrast, the full attention framework operates on all tokens of the video, allowing correlations across both temporal and spatial dimensions simultaneously. To focus on frame correlations, we group and average the tokens in the full attention framework by the frames before conducting further analysis. In Figure 17 and Figure 18, we use wang2024lavie and kong2024hunyuanvideo as representatives of different architectures and present the attention maps across different layers and timesteps in both qualitative and quantitative manners.
@@
Additionally, we present the inter-shot/intra-shot attention probability ratio for different models and layers, as well as the Pearson correlation between this ratio and the distance from transition points (i.e., whether the ratio increases as partition points approach true transitions), with results recorded in Table 6. It can be observed that the inter-shot/intra-shot ratio tends to increase as partition points approach the transition points across different models, further supporting the general presence of the block-diagonal pattern in diffusion models.
Finally, we investigate the potential cause of this pattern by comparing two cases with identical conditions but different seeds: one exhibiting a shot transition and the other only a camera movement (Figure 19). Attention map analysis in Table 7 shows that the case with a shot transition exhibits a stronger block-diagonal pattern, further proving that this attention pattern is strongly correlated with the diffusion model’s understanding of shot transitions.
Figure 19: Comparison of attention maps between two cases with identical conditions but different seeds: one exhibiting a shot transition and the other only showing a camera movement.
Table 6: Inter-shot/Intra-shot attention probability ratio and Pearson correlation for different models and layers.
Model Intra-shot/Inter-shot attention probability ratio Pearson correlation
Wan2.2 (front layer 33%) 5.6792 0.9267,p<<0.0001
Wan2.2 (middle layer 33%) 7.4204 0.9233,p<<0.0001
Wan2.2 (late layer 33%) 6.9483 0.8496,p<<0.0001
Hunyuan (front layer 33%) 4.8514 0.6571,p<<0.0001
Hunyuan (middle layer 33%) 3.4478 0.7507,p<<0.0001
Hunyuan (late layer 33%) 1.5194 0.4969,p<<0.0001
Wan2.1 (front layer 33%) 8.3925 0.7393,p<<0.0001
Wan2.1 (middle layer 33%) 14.1942 0.7319,p<<0.0001
Wan2.1 (late layer 33%) 14.6686 0.6159,p<<0.0001
Table 7: Inter-shot/Intra-shot attention probability ratio and Pearson correlation for two cases (with and without transition).
Model Intra-shot/Inter-shot attention probability ratio Pearson correlation
Case with shot transition
Wan2.1 (front layer 33%) 8.3925 0.7393,p<<0.0001
Wan2.1 (middle layer 33%) 14.1942 0.7319,p<<0.0001
Wan2.1 (late layer 33%) 14.6686 0.6159,p<<0.0001
Case without shot transition
Wan2.1 (front layer 33%) 6.7575 0.3462,p<<0.001
Wan2.1 (middle layer 33%) 4.9148 0.3378,p<<0.001
Wan2.1 (late layer 33%) 6.7575 0.3462,p<<0.001
@@ D.2 Impact of Mask Layers on Results
To validate the generalizability of our heuristic mask-layer selection, we apply the same approach to other architectures: masking the later half of layers in U-Net and the middle half of layers in DiT. We use VideoCrafter2 chen2024videocrafter2 and Wan2.2 wan2025wan as base models, with results presented in Table 8. Due to VideoCrafter2’s relatively limited duration and representational capacity, the overall Transition Control Scores are relatively low. However, masking later layers still leads to clearer shot-transition behavior. Figure 21 shows that selecting the appropriate masking layer can indeed improve multi-shot video generation performance.
Table 8: Quantitivate results for masking different layers in new different frameworks. The best are in bold.
Semantic Visual Subject↑ Background↑
Score↑ Gap↓ Score↑ Gap↓
Ablation for VideoCrafter2 (Unet)
Mask front 1/2 layers 0.0889 0.8313 0.3630 0.8338 0.3603 0.9594 0.9723 0.6287 0.2205
Mask late 1/2 layers 0.1889 0.7739 0.1980 0.7766 0.3416 0.9728 0.9693 0.6258 0.2213
Ablation for Wan2.2 (DiT)
Mask front 1/2 layers 0.3811 0.9425 0.5976 0.9380 0.5205 0.9570 0.9624 0.6298 0.2073
Mask middle 1/2 layers 0.5982 0.6752 0.2084 0.7234 0.1458 0.9474 0.9636 0.6395 0.2140
Mask late 1/2 layers 0.0085 - - - - 0.9466 0.9646 0.6305 0.1943
Figure 21: Results of masking different layers in new different frameworks.
@@ D.3 Qualitative Evaluation for Ablation Studies
D.4 User Study
To complement the experimental results, we report the findings of our user study in this section. Specifically, we select 20 prompts for user evaluation, where participants rate each result on a scale from 1 to 5. We evaluate different methods from the perspective of Transition Control and Overall Consistency, with the MOS results presented in Table 9. It can be observed that CineTrans also demonstrates strong performance in terms of user preference.
Table 9: Results of User Study
Model Transition Control Consistency
CineTrans-DiT (ours) 4.60 $\pm$ 0.50 4.15 $\pm$ 0.75
CineTrans-Unet (ours) 4.75 $\pm$ 0.44 4.10 $\pm$ 0.72
StoryDiffusion+CogVideoXI2V - 3.95 $\pm$ 0.76
HunyuanVideo+Cinematron 3.60 $\pm$ 1.95 3.80 $\pm$ 0.68
HunyuanVideo 3.45 $\pm$ 1.88 3.50 $\pm$ 0.76
CogVideoX 2.50 $\pm$ 1.24 3.05 $\pm$ 0.60
Wanx2.1-T2V-turbo 3.25 $\pm$ 1.77 3.45 $\pm$ 0.69
D.5 Achieving Smoother Transitions through Soft Masking
This section investigates soft masking strategies applied to the hard mask mechanism proposed in 4.3, with the aim of assessing whether soft masking can enable smoother shot transitions or higher consistency. Specifically, we explore strategies including time-dependent penalty and timestep-dependent penalty. The following will present the details of these strategies along with the corresponding experimental results.
Time-Dependent Penalty. In the hard masking scheme, the mask matrix is initially set to be fully invisible, allowing tokens within each shot to interact for multi-shot video generation. The time-dependent penalty allows token interactions near shot boundaries. Specifically, we initialize the mask matrix using a Gaussian decay based on the distance between frames:
$\mathcal{M}_{ij}=e^{-\dfrac{(i-j)^{2}}{2\sigma^{2}}}.$ (8)
This is then mapped to the range $0,-\infty)$, with all tokens within each shot being fully visible subsequently, thus forming a soft masking strategy around shot boundaries. The smoothness of the transition is controlled by the parameter $\sigma$. The evaluation results are presented in Table 10 and Figure 23, where $L$ denotes the sequence length. As the soft mask approaches full visibility, the transition control effect weakens until no shot transition occurs. With the appropriate hyperparameter settings (e.g., $\sigma=L/12$), the visual break at the transition boundary becomes less obvious. However, excessively large $\sigma$ values lead to minimal difference across shot compositions, reducing the content diversity in the multi-shot video and resulting in a more uniform appearance, which explains the decline in semantic consistency.
Figure 23: An illustration of the Time-Dependent Penalty.
Table 10: Quantitative results for Time-Dependent Penalty. The best are in bold. Smaller $\sigma$ values approach the hard mask.
Semantic Visual Subject↑ Background↑
Score↑ Gap↓ Score↑ Gap↓
Ablation for CineTrans-DiT (training-free)
$\sigma=L/4$ 0.0220 0.8685 0.4261 0.8231 0.2499 0.9629 0.9750 0.6513 0.2025
$\sigma=L/6$ 0.0700 0.8238 0.2411 0.8248 0.2431 0.9445 0.9606 0.6514 0.2073
$\sigma=L/12$ 0.4036 0.7987 0.2005 0.8186 0.2139 0.9400 0.9544 0.6562 0.2086
Hard Mask 0.6564 0.7838 0.1772 0.7844 0.1943 0.9618 0.9746 0.6556 0.2093
Ablation for CineTrans-DiT (trained)
$\sigma=L/4$ 0 - - - - 0.9541 0.9660 0.6453 0.2004
$\sigma=L/6$ 0.0455 0.7944 0.1885 0.8377 0.3074 0.9424 0.9598 0.6455 0.2105
$\sigma=L/12$ 0.5193 0.7902 0.1796 0.8145 0.2575 0.9519 0.9659 0.6489 0.2095
Hard Mask 0.7003 0.7858 0.1552 0.7874 0.1901 0.9673 0.9775 0.6508 0.2109
Timestep-Dependent Penalty. In contrast, the Diffusion-Timestep-Dependent Penalty focuses on the timestep of the denoising process, making the invisible region of the mask partially visible during the early denoising steps and gradually approaching full invisibility as the process progresses. Compared to the Time-Dependent Penalty, which focuses on shot boundaries, this soft masking enhances interactions between all inter-shot tokens. As shown in Figure 24, the increased token correlations result in a significant reduction in compositional differences between shots, leaving only a residual boundary transition effect. This comes at the cost of multi-shot content diversity, as reflected by the substantial drop in the Transition Control Score in Table 11. Therefore, the Diffusion-Timestep-Dependent Penalty tends to strengthen inter-shot consistency, while its effect on smoother transitions is less pronounced than that of the time-dependent penalty.
Table 11: Quantitative results for Timestep-Dependent Penalty. The best are in bold. Smaller $t$ values approach the hard mask.
Method Transition Control Score Inter-shot Consistency
Semantic Visual
Score↑ Gap↓ Score↑ Gap↓
Ablation for CineTrans-DiT (training-free)
$t=1.0$ 0 - - - -
$t=0.5$ 0 - - - -
Hard Mask 0.6564 0.7838 0.1772 0.7844 0.1943
Ablation for CineTrans-DiT (trained)
$t=1.0$ 0 - - - -
$t=0.5$ 0.0741 0.7864 0.2211 0.8029 0.3655
Hard Mask 0.7003 0.7858 0.1552 0.7874 0.1901
Figure 24: An illustration of the Timestep-Dependent Penalty.
D.6 Overlapping Ambiguous Shot Boundaries
The results in Table 1 demonstrate that our model maintains precise control over shot transitions at specified transition points using the mask mechanism. In this section, we examine the model’s performance under overlapping and ambiguous shot boundaries. Specifically, we introduce fluctuations of ±$t$ frames (temporal slices) around the fixed transition points during inference, which generates overlapping and ambiguous boundaries in the denoising process. Quantitative results are shown in Table 12 and Figure 25. It can be observed that, even under overlapping and ambiguous shot boundaries, the model retains a reasonable Transition Control Score and produces effects resembling fade-in/fade-out transitions to some extent. These results confirm the robustness of the mask mechanism, demonstrating that the model retains effective control even in the presence of ambiguous or overlapping shot boundaries.
Figure 25: A Result of Overlapping Ambiguous Shot Boundary.
Table 12: Quantitative results for Overlapping Ambiguous Shot Boundaries. The best are in bold. Smaller $t$ values approach the fixed transition points.
Method Transition Control Score Inter-shot Consistency
Semantic Visual
Score↑ Gap↓ Score↑ Gap↓
Ablation for CineTrans-DiT (training-free)
No Mask 0.2051 0.5924 0.3421 0.5274 0.3574
$t=2$ 0.2698 0.7227 0.1918 0.8109 0.2612
$t=1$ 0.4123 0.7344 0.1846 0.8044 0.2299
Fixed Transition Points 0.6564 0.7838 0.1772 0.7844 0.1943
Ablation for CineTrans-DiT (trained)
No Mask 0.2112 0.6532 0.3422 0.6087 0.3312
$t=2$ 0.4240 0.7526 0.1976 0.8311 0.3082
$t=1$ 0.5461 0.7302 0.1837 0.8245 0.2764
Fixed Transition Points 0.7003 0.7858 0.1552 0.7874 0.1901